OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of three-fifths?

OpenStudy (anonymous):

y-b=m(x-a) a= 3 b = -2 m =3/5

OpenStudy (anonymous):

y-(-2)=3/5(x-3) =5(y+2)=3(x-3) =5y+10=3x-15 <=>3x-5y-25=0 or equivolently 3/5x-y-5=0

OpenStudy (anonymous):

y=mx+b which means, nevermind.. lol I see they have it.

OpenStudy (anonymous):

-2=3/5(3) +b

OpenStudy (anonymous):

Ullere thank u for helping but where did you get the 5 from?

OpenStudy (anonymous):

Which 5?

OpenStudy (anonymous):

3/5 is what u gave as a slope

OpenStudy (anonymous):

\[y-(-2)=\frac{3}{5}(x-3)\]

OpenStudy (anonymous):

ooh never mind you divided both sides by 5 to get the 5 in the second line?

OpenStudy (anonymous):

\[y+2=\frac{3}{5}x-\frac{9}{5}\]

OpenStudy (anonymous):

\[5y+10=3x-9\]\[3x-5y-19=0\]

OpenStudy (anonymous):

Yeah, you said standard form and sometimes those questions are snippy and it being x-y+c=0 or y=mx+c etc. They are the same either way.

OpenStudy (anonymous):

Woah sarah is right. I made 3*3 = 15, silly mistake.

OpenStudy (anonymous):

y-(-2)=3/5(x-3) =5(y+2)=3(x-3) =5y+10=3x-9 <=>3x-5y-19=0 or equivolently 3/5x-y-19/5=0